# Look as You Leap: Planning Simultaneous Motion and Perception for High-d o f Robots

### v1

@@ V-C Human Detection in a Nursing Setting
Table II summarizes the average performance over all runs, reporting metrics such as planning time, path length, face detection rate (for detecting at least one nurse or both nurses), and average face detection confidence score. We observe that Manifold-line tends to choose a shorter path by moving behind the nurses, as shown in Fig. 7, which often leads to poor visibility of their faces. In contrast, Neural-PS-PRM-line consistently favors slightly longer and cluttered paths that provide better visibility of the nurses’ faces.
As shown in Table II, Neural-PS-PRM-line significantly outperforms Manifold-line in both face detection rate and average confidence score. This improvement comes at the cost of increased path length and planning time. Specifically, Neural-PS-PRM-line achieves over 98% success in detecting at least one nurse and over 63% for detecting both nurses, while the baseline method fails entirely in the two-face detection task. An example of the routes selected by our method and the baseline is shown in Fig. 7.
In summary, these experiments confirm that our approach PS-PRM is effective across different environments and is a suitable candidate for planning simultaneous perception and motion for high-DoF robots in practical human-robot interaction applications.
@@
Motion Planning Human Detection
@@
TABLE II: Comparison of Manifold-line and Neural-PS-PRM-line methods in a simulated nursing environment. “Plan. Time” refers to the average planning time, with Neural-PS-PRM-line also accounting for the time required to construct the PRM. “Path Len.” represents the average path length. “Det. Rate (1 face)” and “Det. Rate (2 faces)” indicate the face detection rates for one or both nurses, respectively. “Conf. Score” refers to the average confidence score of face detections for the two nurses.
V-D Planning in a Dynamic Environment
As shown in Fig. 8, we construct a dynamic environment containing a human, a cabinet, a shelf, and other common household objects such as a sofa, table, and chair. The robot is tasked with planning a path from its initial position to reach one of two goal locations (the cabinet or the shelf), while continuously detecting key facial features of the human. This experiment focuses on a face detection and tracking task, utilizing DeepFace and RetinaFace to assess real-time face detection performance. All trials are conducted using the Stretch 2 robot. During execution, the human moves and rotates from side to side, requiring the robot to dynamically select which goal to pursue in order to maintain high face detection performance.
For the specific scenario illustrated in Fig. 8, we evaluate both Manifold-line and Neural-PS-PRM-line over 100 independent runs. Each Manifold-line trial is given a maximum of 60 seconds to compute a plan. In contrast, Neural-PS-PRM-line constructs a PRM and assigns perception-aware scores using a trained neural surrogate model based on face detection confidence. Unlike Manifold-line, which does not account for the human’s dynamic movements, Neural-PS-PRM-line supports online replanning during execution as described in Sec. IV-C. Table III presents the average results across all runs, evaluating metrics such as planning time, path length, face detection rate, and average confidence score.
As expected, Manifold-line consistently selects a single goal and results in very low face detection performance. In contrast, Neural-PS-PRM-line initially chooses the path toward the shelf on the left, as the human begins facing that direction. During execution, the robot replans in real time and switches to the path toward the cabinet on the right as the human turns, thereby maintaining better visibility of the face. As shown in Table III, Neural-PS-PRM-line significantly outperforms Manifold-line in both face detection rate and average confidence score, while also demonstrating successful online replanning. An example of the routes selected by our method is shown in Fig. 8.
@@ V-D Planning in a Dynamic Environment
TABLE III: Comparison of Manifold-line and Neural-PS-PRM-line methods in a dynamic, human-aware navigation scenario. ”Plan. Time” refers to the average time spent planning, with Neural-PS-PRM-line including PRM construction. ”#Plans” indicates the average number of replanning instances during execution. ”Path Len.” is the average path length in radians. ”Det. Rate” is the percentage of frames in which the person’s face was successfully detected. ”Conf. Score” represents the average confidence score of detected faces.
Figure 9: In this real robot experiment, the robot navigates from the start to goal while trying to detect a cup on the table. The figure shows the motion paths for two methods: Manifold-line (orange) and PS-PRM-line (blue). Camera views for each method are shown progressively from top left to bottom right.
V-E Mobile Manipulator Planning in Blind Spot Alley
@@
V-F Real Robot Experiments
V-F1 Static Object Tracking Task
@@
V-F2 Dynamic Human Detection Task
@@
In this experiment, illustrated in Fig. 10, we demonstrate the deployment of our system in a realistic nurse training scenario. The environment includes a nurse engaged in her workspace, surrounded by typical elements such as work tables, a chair, a platform, a garbage can, and two cabinets that serve as potential goal locations where the robot needs to grab some tools for the nurse. The task requires the Stretch 2 robot to navigate to one of these locations while maintaining reliable face detection of the nurse As the nurse rotates during the task, the robot must adapt to changes in facial visibility. Our system initially plans a path to the left cabinet (Goal 1) when the nurse faces left, but dynamically replans to the right cabinet (Goal 2) as she turns right, preserving optimal visibility. This experiment highlights how Neural-PS-PRM-line enables perception-aware, online replanning to support continuous coordination between motion ad perception, a critical requirement in nurse training and collaborative medical environments.
@@
In summary, these real-robot experimenta show that PS-PRM is suitable for task requiring simultaneous perception and motion in dynamic environments.
@@ VI Discussion
This direction opens several promising avenues for future research. First, while the neural surrogate model significantly reduces computation time, it is currently trained offline under fixed environmental conditions. Future work could explore domain adaptation techniques or continual learning strategies to generalize across environments and sensor configurations. Second, incorporating uncertainty estimates from the surrogate model may enhance robustness, particularly in ambiguous or noisy scenes. Third, learning joint representations that fuse geometry, semantics, and visibility could yield tighter integration between perception and planning, especially in occlusion-heavy settings.

### v2

@@ V-D Human Tracking in a Nursing Setting
Table IV summarizes the average performance over all runs, reporting metrics such as planning time, path length, face detection rate (for detecting at least one nurse or both nurses), and average face detection confidence score. We observe that Manifold-line tends to choose a shorter path by moving behind the nurses, as shown in Fig. 7, which often leads to poor visibility of their faces. In contrast, Neural-PS-PRM-line consistently favors slightly longer and cluttered paths that provide better visibility of the nurses’ faces.
As shown in Table IV, Neural-PS-PRM-line significantly outperforms Manifold-line in both face detection rate and average confidence score. This improvement comes at the cost of increased path length and planning time. Specifically, Neural-PS-PRM-line achieves over 98% success in detecting at least one nurse and over 63% for detecting both nurses, while the baseline method fails entirely in the two-face detection task. An example of the routes selected by our method and the baseline is shown in Fig. 7.
Fig. 8 illustrates a scenario with conflicting target placements, where the planner must balance multiple perception objectives. In this experiment, the two nurses are spatially separated such that maintaining high-quality perception of both targets simultaneously is not feasible. To resolve this conflict, the planner does distance-based target-selection. As the Stretch robot moves from left to right, it initially prioritizes tracking Nurse 1, and shifts attention toward Nurse 2 as the robot approaches closer, reflecting the changing importance of the perception objectives. Given the requirement of constraint satisfaction for this problem, the constraint tolerance is increased for this experiment to work, allowing the camera to aim away from one object into another during planning.
@@
Motion Planning Human Tracking
@@
TABLE IV: Comparison of Manifold-line and Neural-PS-PRM-line methods in a simulated nursing environment. “Plan. Time” refers to the average planning time, with Neural-PS-PRM-line also accounting for the time required to construct the PRM. “Path Len.” represents the average path length. “Det. Rate (1 face)” and “Det. Rate (2 faces)” indicate the face detection rates for one or both nurses, respectively. “Conf. Score” refers to the average confidence score of face detections for the two nurses.
Figure 8: This simulation experiment illustrates how our planner handles conflicting perception objectives using distance-based priority. In this scenario, the two nurses are spatially separated. As the Stretch robot moves from left to right, it initially prioritizes tracking Nurse 1 and shifts attention to Nurse 2.
V-E Planning in a Dynamic Environment
As shown in Fig. 9, we construct a dynamic environment containing a human, a cabinet, a shelf, and other common household objects such as a sofa, table, and chair. The robot is tasked with planning a path from its initial position to reach one of two goal locations (the cabinet or the shelf), while continuously detecting key facial features of the human. This experiment focuses on a face detection and tracking task, utilizing RetinaFace [21] to assess real-time face detection performance. All trials are conducted using the Stretch 2 robot. During execution, the human moves and rotates from side to side, requiring the robot to dynamically select which goal to pursue in order to maintain high face detection performance.
For the specific scenario illustrated in Fig. 9, we evaluate both Manifold-line and Neural-PS-PRM-line over 100 independent runs. Each Manifold-line trial is given a maximum of 60 seconds to compute a plan. In contrast, Neural-PS-PRM-line constructs a PRM and assigns perception-aware scores using a trained neural surrogate model based on face detection confidence. Unlike Manifold-line, which does not account for the human’s dynamic movements, Neural-PS-PRM-line supports online replanning during execution as described in Sec. IV-C. Table V presents the average results across all runs, evaluating metrics such as planning time, path length, face detection rate, and average confidence score.
As expected, Manifold-line consistently selects a single goal and results in very low face detection performance. In contrast, Neural-PS-PRM-line initially chooses the path toward the shelf on the left, as the human begins facing that direction. During execution, the robot replans in real time and switches to the path toward the cabinet on the right as the human turns, thereby maintaining better visibility of the face. As shown in Table V, Neural-PS-PRM-line significantly outperforms Manifold-line in both face detection rate and average confidence score, while also demonstrating successful online replanning. An example of the routes selected by our method is shown in Fig. 9.
@@ V-E Planning in a Dynamic Environment
TABLE V: Comparison of Manifold-line and Neural-PS-PRM-line methods in a dynamic, human-aware scenario. ”Plan. Time” refers to the average time spent planning, with Neural-PS-PRM-line including PRM construction. ”#Plans” indicates the average number of replanning instances made during execution. ”Path Len.” is the average path length in radians. ”Det. Rate” is the percentage of frames in which the person’s face was successfully detected. ”Conf. Score” represents the average confidence score of the detected faces.
Figure 10: In this real robot experiment, the robot plans a trajectory from the start to goal while trying to detect a cup on the table. The figure shows the motion paths for two methods: Manifold-line (orange) and PS-PRM-line (blue). Camera views for each method are shown progressively from top to bottom, left to right.
V-F Mobile Manipulator Planning in Blind Spot Alley
@@
V-G Real Robot Experiments
V-G1 Static Object Tracking Task
@@
Figure 12: This real-robot experiment demonstrates dynamic human (perception target), obstacle and goal movement. The Stretch 2 robot (highlighted in blue) must reach the goal cabinet while maintaining reliable face detection. As the person rotates (B), the robot switches between right- and left-side paths to preserve visibility. The obstacle cabinet is then moved (C), prompting collision-avoidance replanning, followed by a final replan as the goal cabinet changes position (D).
Figure 13: This real-robot experiment demonstrates perception-constrained motion planning for human activity recognition on a 6-DoF UR5 manipulator. The left trajectory is generated without perception costs and follows a short path to the goal, during which the wrist-mounted camera view of the person is largely occluded, resulting in missed human activity observations. In contrast, the PS-PRM–generated trajectory takes a longer detour to maintain visibility of the person, enabling sustained human activity recognition throughout most parts of the execution. Trajectories are shown as dotted lines, colored red when the person is occluded and green when the person is visible.
V-G2 Dynamic Human Tracking Task
@@
In this experiment, illustrated in Fig. 11, we demonstrate the deployment of our system in a realistic nurse training scenario. The environment includes a nurse engaged in her workspace, surrounded by typical elements such as work tables, a chair, a platform, a garbage can, and two cabinets that serve as potential goal locations where the robot needs to grab some tools for the nurse. The task requires the Stretch 2 robot to navigate to one of these locations while maintaining reliable face detection of the nurse. As the nurse rotates during the task, the robot must adapt to changes in facial visibility. Our system initially plans a path to the left cabinet (Goal 1) when the nurse faces left, but dynamically replans to the right cabinet (Goal 2) as she turns right, preserving optimal visibility. This experiment highlights how Neural-PS-PRM-line enables evaluating trajectories on real-time and replanning, a critical requirement in nurse training and collaborative medical environments.
V-G3 Human Tracking with Moving Obstacle and Goal
In this experiment (illustrated in Fig. 12), we demonstrate our system’s performance in a realistic, dynamic scenario. The environment contains a seated person reading a book, surrounded by two cabinets that serve respectively as an obstacle and a goal location the robot must reach. The task requires the Stretch 2 robot to reach a configuration near the goal cabinet while maintaining reliable face tracking of the person. As the person naturally rotates during the task, the robot must adapt to changes in their pose and the obstacle locations. When the person initially faces right, the robot selects a path on the right side; as the person turns left, the system dynamically replans and shifts to a left-side path to preserve optimal perception. We then introduce additional environmental changes: the obstacle cabinet is moved mid-execution, forcing the robot to replan in real time to avoid collision, and we further demonstrate that the robot can plan toward a moving goal as the goal cabinet changes position. For experiment setup, we use an OptiTrack motion capture system to track the positions and orientations of both the cabinets and the human; this could be replaced with alternative pose tracking solutions.
By exploiting the biased sampler and triggering replanning via movement thresholds (determined by collision paddings), the planner achieves execution rates of up to 2Hz. It also generated an average of 8 trajectories per run, more than the 3 environment changes. This highlights the capability to plan multiple times on single scene alterations. Overall, this experiment shows how Neural-PS-PRM-line enables collision- and perception-aware online replanning, supporting continuous coordination between motion and perception in the presence of moving obstacles and shifting perception target and motion goal locations.
V-G4 Manipulator Grasping Task with Human Tracking Constraints
In this experiment (illustrated in Fig. 13), we demonstrate our system’s performance on a real-robot grasping task using a UR5 manipulator. The environment contains a seated person reading a book, and the robot is required to grasp an object while continuously monitoring the person’s activity. A wooden board introduces additional occlusion that can partially block the wrist-mounted camera view. The robot’s task is to move from left to right while maintaining human activity detection for as long as possible.
During the execution, the person alternates between reading a book, picking up a water bottle to drink, and then continuing to read. Human activity recognition is performed using a vision–language–based activity detection model [42]. As shown in the left subfigure, the baseline method PS-PRM-line, which does not incorporate occlusion reasoning, selects the shortest path to the goal. However, this trajectory fails to observe the person’s activity for most of the execution and, in particular, misses the drinking action. In contrast, as shown in the right subfigure, Neural-PS-PRM-line plans a longer trajectory that moves above the wooden board, maintaining visibility of the person throughout most of the motion. As a result, it successfully captures the drinking activity and accurately detects the time intervals during which the person is reading.
This real-robot experiment demonstrates the practical applicability of our method to manipulation tasks that require human awareness under occlusion.
@@ V-G4 Manipulator Grasping Task with Human Tracking Constraints
In summary, these real-robot experiments show that PS-PRM is suitable for tasks requiring simultaneous perception and motion in dynamic environments.
@@ VI Discussion
This direction opens several promising avenues for future research. First, while the neural surrogate model significantly reduces computation time and is robot- and environment-agnostic, it is currently trained offline for a fixed set of objects and under a limited set of environmental settings, such as light conditions. Future work could explore domain adaptation techniques or continual learning strategies to generalize across environments, sensor configurations or object types. Second, incorporating uncertainty estimates from the surrogate model may enhance robustness, particularly in ambiguous or noisy scenes. Third, there remains opportunities for further GPU parallelization, such as incorporating ideas from [2] to further accelerate PRM construction and query. An interesting extension of this work is to explore how the framework can be adapted to a broader range of perception tasks and application domains, such as semantic segmentation or 3D reconstruction.